A good is represented by a market demand curve Q = 100 – P. In this market, there are an unlimited number of potential firms whose cost curve is given as TC = Q + Q2.

 

a)    What is the long run equilibrium price, assuming free entry of firms?

(8 marks)

 

b)    How many firms will there be?



1
Expert's answer
2021-07-14T12:01:06-0400

a) Long run equilibrium occurs when Marginal cost = Average total cost = P

Total cost = "q+q^2"

MC=1 + 2q

ATC = 1+ q

1+2q=1+q

q=0

P=1

So, market demant Q=100-P = 100-1=99

b)

"nq=Q \\

n = \frac{Q}{q} \\

n = \frac{99}{0} = infinity"

This technically means that there is an unlimited number of firms in this market.